# Early Stages of Self-Healing at Tungsten Grain Boundaries from Ab Initio Machine Learning Simulations

### v1

@@ 4.1 MLIP-W development
In the exploration step, an iterative active learning protocol was implemented to refine the potential. Using the models described above, short MD simulations (on the order of a few tens of picoseconds) in the NVE ensemble were performed at temperatures of 300 K, 650 K, and 1000 K, corresponding to the three different regimes identified by Beyerlein et al. [13], each with a time step of 5 femtoseconds and 10,000 MD steps. In each cycle, 10 Frenkel pairs were introduced with a separation of $\geq 3$, allowing for natural defect evolution near GBs. The aforementioned potential ensemble was then used to measure the maximum standard deviation of the predicted atomic forces of the samples obtained by MD, i.e.,
@@
where $\epsilon_{t}$ and $R_{t}$ denote the maximum force deviation and the atomic coordinates of a given structure $t$, respectively. The index $i$ runs over all atoms in the structure and the index $w$ denotes the $w$th potential model in the ensemble, meaning that the atomic forces $F_{w,i}$ are ensemble averaged. Four potentials are often considered sufficient to construct this ensemble. The above equation was then used as a criterion to select candidate structures from the MD trajectories, i.e., those structures whose $\epsilon_{t}$ satisfied the inequality $\sigma_{f}^{\text{min}}<\epsilon_{t}<\sigma_{f}^{\text{max}}$ were selected and sent to the next step. In practice, $\sigma_{f}^{\text{min}}$ and $\sigma_{f}^{\text{max}}$ are set empirically, with no fixed rule, just to ensure that a reasonable fraction of candidates are selected in each iteration. In this work, $\sigma_{f}^{\text{min}}=0.20$ and $\sigma_{f}^{\text{max}}=1.00$ were used.
@@ 4.2 DFT calculation details
Two representative systems were fully relaxed using DFT: (i) a 5 × 5 × 5 supercell of bulk W, containing 250 atoms, to capture the intrinsic properties of the BCC lattice; and (ii) the W110delimited-⟨⟩110$\langle$ 110$\rangle$ 110 ⟩/W112delimited-⟨⟩112$\langle$ 112$\rangle$ 112 ⟩ interface, with 456 atoms in a cell of dimensions 31.09 Å × 10.98 Å × 31.72 Å. The construction of the W110delimited-⟨⟩110$\langle$ 110$\rangle$ 110 ⟩/W112delimited-⟨⟩112$\langle$ 112$\rangle$ 112 ⟩ interface follows the methodology described in Ref. [76], where two 6-layer slabs are joined along the $z$-axis. The (110) surface was rotated by 55 relative to the (112) surface, and a small expansion ($\approx 1\%$) is applied along the $x$-direction to accommodate the non-coherent nature of the interface. A plane-wave cutoff energy of 400 eV was used in all calculations, and structural relaxations continued until the residual Hellmann–Feynman forces were below 0.03 eV Å-1. For the bulk system, a 3 × 3 × 3 Monkhorst-Pack k-point mesh was used to ensure adequate Brillouin zone sampling. For the larger W110delimited-⟨⟩110$\langle$ 110$\rangle$ 110 ⟩/W112delimited-⟨⟩112$\langle$ 112$\rangle$ 112 ⟩ interface cell, only the Gamma point was selected, which has been shown to provide sufficient accuracy in a previous work [77]. Following these relaxations, the resulting equilibrium lattice constant for W was found to be 3.172 Å, in good agreement with the experimental value of 3.165 Å [78].
@@ 4.3 MD simulations
All MD simulations were performed using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [79] patched with the DeePMD-kit [42]. Alongside our newly developed DeePMD model for W, two widely used empirical IPs were employed for comparison: the EAM potential developed by Marinica et al. [50], commonly applied to describe tungsten–tungsten interactions in irradiation damage studies, and the Tersoff potential parameterized by Juslin and Wirth [80], originally developed for covalent systems which incorporate a bond-order framework that captures short-range angular interactions and enables the modeling of complex defect structures in W. For system minimization, periodic boundary conditions (PBC) were imposed in all directions, effectively simulating an infinite crystal. A 0.5 fs timestep was used to ensure numerical stability under defect introduction, and a conjugate gradient (CG) relaxation with zero energy and force tolerances was employed, allowing up to 10,000 steps if convergence was not reached earlier. An anisotropic relaxation scheme limited the maximum fractional deformation of the simulation box to 0.001 per minimization step.
@@ 4.4 FaVAD
$=$ $\sum_{nlm}^{NLM}c^{(i)}_{nlm}g_{n}(r)Y_{lm}\left(\hat{r}\right),$ (4)
where $\vec{r}_{ij}$ is the internuclear distance difference vector between the atoms $i$ and $j$; $\sigma^{2}_{\textrm{atom}}$ is associated to the broadening of the atomic position and $f_{\textrm{cut}}\left(|\vec{r}_{ij}|\right)$ is a smooth cutoff function which removes atoms $j$ with a distance $|\vec{r}_{ij}|$ exceeding $r^{c}_{ij}$ from the local atomic density computation around the $i$-th atom. Here, $r^{c}_{ij}$ denotes the cutoff radius used in the calculation of the local atomic density around the $i$-th atom. $Y_{lm}(\hat{r})$ are the spherical harmonics and $\hat{r}$ is a unit vector in the $\vec{r}$ direction [39]. Finally, the expansion coefficients were calculated by an orthonormal set of radial functions $g_{n}(r)$ as $c^{(i)}_{nlm}=\langle g_{n}Y_{lm}|\rho_{i}\rangle$ [82]. The sum over the order $m$ of the squared modulus of the coefficients $c_{nlm}$ is invariant under rotations around the central atom that defines the DV as a fingerprint of the particular local atomic environment as

### v2

@@ 3 Discussion
Moreover, our descriptor-based analysis reveals that configurations generated with MLIP-W distinctly differentiate between defect-related environments and thermal fluctuations. In contrast, configurations obtained with traditional IPs, particularly the Tersoff model, exhibit substantial overlap between defect-induced and thermally driven atomic displacements. This is consistent with the blurred separation observed in the per-atom potential energy analysis. Finally, an Arrhenius analysis of defect annihilation rates yields a migration energy of 0.048 eV when using the MLIP-W, in agreement with the reported value range [62, 76].
@@ 4.1 MLIP-W development
In the exploration step, an iterative active learning protocol was implemented to refine the potential. Using the models described above, short MD simulations (on the order of a few tens of picoseconds) in the NVE ensemble were performed at temperatures of 300 K, 650 K, and 1000 K, corresponding to the three different regimes identified by Beyerlein et al. [13], each with a time step of 5 femtoseconds and 10,000 MD steps. In each cycle, 10 Frenkel pairs were introduced with a separation of $\geq 3$ Å, allowing for natural defect evolution near GBs. The aforementioned potential ensemble was then used to measure the maximum standard deviation of the predicted atomic forces of the samples obtained by MD, i.e.,
@@
where $\epsilon_{t}$ and $R_{t}$ denote the maximum force deviation and the atomic coordinates of a given structure $t$, respectively. The index $i$ runs over all atoms in the structure and the index $w$ denotes the $w$th potential model in the ensemble, meaning that the atomic forces $F_{w,i}$ are ensemble averaged. Four potentials are often considered sufficient to construct this ensemble. The above equation was then used as a criterion to select candidate structures from the MD trajectories, i.e., those structures whose $\epsilon_{t}$ satisfied the inequality $\sigma_{f}^{\text{min}}<\epsilon_{t}<\sigma_{f}^{\text{max}}$ were selected and sent to the next step. In practice, $\sigma_{f}^{\text{min}}$ and $\sigma_{f}^{\text{max}}$ are set empirically, with no fixed rule, just to ensure that a reasonable fraction of candidates is selected in each iteration. In this work, $\sigma_{f}^{\text{min}}=0.20$ and $\sigma_{f}^{\text{max}}=1.00$ were used.
@@ 4.2 DFT calculation details
Two representative systems were fully relaxed using DFT: (i) a 5 × 5 × 5 supercell of bulk W, containing 250 atoms, to capture the intrinsic properties of the BCC lattice; and (ii) the $\langle 110\rangle/W$$\langle 112\rangle$ interface, with 456 atoms in a cell of dimensions 31.09 Å × 10.98 Å × 31.72 Å. The construction of the $\langle 110\rangle/W$$\langle 112\rangle$ interface follows the methodology described in Ref. [82], where two 6-layer slabs are joined along the $z$-axis. The (110) surface was rotated by 55 relative to the (112) surface, and a small expansion ($\approx 1\%$) is applied along the $x$-direction to accommodate the non-coherent nature of the interface. A plane-wave cutoff energy of 400 eV was used in all calculations, and structural relaxations continued until the residual Hellmann–Feynman forces were below 0.03 eV Å-1. For the bulk system, a 3 × 3 × 3 Monkhorst-Pack k-point mesh was used to ensure adequate Brillouin zone sampling. For the larger $\langle 110\rangle/W$$\langle 112\rangle$ interface cell, only the Gamma point was selected, which has been shown to provide sufficient accuracy in a previous work [83]. Following these relaxations, the resulting equilibrium lattice constant for W was found to be 3.172 Å, in good agreement with the experimental value of 3.165 Å [84].
@@ 4.3 MD simulations
All MD simulations were performed using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [85] patched with the DeePMD-kit [43]. Alongside the DeePMD model for W that we developed, two widely used empirical IPs were employed for comparison: the EAM potential developed by Marinica et al. [51], commonly applied to describe tungsten–tungsten interactions in irradiation damage studies, and the Tersoff potential parameterized by Juslin and Wirth [86], originally developed for covalent systems which incorporate a bond-order framework that captures short-range angular interactions and enables the modeling of complex defect structures in W. For system minimization, periodic boundary conditions (PBC) were imposed in all directions, effectively simulating an infinite crystal. A 0.5 fs timestep was used to ensure numerical stability under defect introduction, and a conjugate gradient (CG) relaxation with zero energy and force tolerances was employed, allowing up to 10,000 steps if convergence was not reached earlier. An anisotropic relaxation scheme limited the maximum fractional deformation of the simulation box to 0.001 per minimization step.
@@ 4.4 FaVAD
[…]
@@
Data Availability
The datasets generated and analyzed during the current study are available from the corresponding author upon reasonable request.
@@ Acknowledgments
Author contributions
J.S.-R., R.I., R.G.-A. and P.M.P. designed the research. J.S.-R. performed the simulations. J.S.-R., P.M.P. and F.J.D.-G. analyzed the data. J.S.-R., P.M.P. and F.J.D.-G. wrote the paper. R.I. and R.G.-A. supervised the project. All authors discussed the results and commented on the manuscript.
Competing interests
The authors declare no competing interests.